OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (-2, 6) and has a slope of -1? a) x + y = -4 b) x + y = 4 c) x - y = -4 d) x - y = 4

OpenStudy (anonymous):

\[y - y_{1} = m(x - x_{1})\]\[y - 6 = -1(x - (-2))\]\[y - 6 = -1(x + 2)\] Can you simplify from here?